### v1

# Continual Learning for Image Captioning through Improved Image-Text Alignment

Abstract
Generating accurate and coherent image captions in a continual learning setting remains a major challenge due to catastrophic forgetting and the difficulty of aligning evolving visual concepts with language over time. In this work, we propose a novel multi-loss framework for continual image captioning that integrates semantic guidance through prompt-based continual learning and contrastive alignment. Built upon a pretrained ViT-GPT-2 backbone, our approach combines standard cross-entropy loss with three additional components: (1) a prompt-based cosine similarity loss that aligns image embeddings with synthetically constructed prompts encoding objects, attributes, and actions; (2) a CLIP-style loss that promotes alignment between image embeddings and target caption embedding; and (3) a language-guided contrastive loss that employs a triplet loss to enhance class-level discriminability between tasks. Notably, our approach introduces no additional overhead at inference time and requires no prompts during caption generation. We find that this approach mitigates catastrophic forgetting, while achieving better semantic caption alignment compared to state-of-the-art methods. The code can be found via the following link https://github.com/Gepardius/Taetz_Bordelius_Continual_ImageCaptioning.
Keywords Image Captioning $\cdot$ Continual Learning $\cdot$ Image-Text Alignment
Figure 1: Overview of the proposed multi-objective training approach combining prompt-based, CLIP-based cosine similarity loss and triplet loss.
1 Introduction
Understanding how to generate natural language descriptions of images is a fundamental problem at the intersection of computer vision and natural language processing. Recent advances in deep learning have enabled remarkable progress in static image captioning [1], where large-scale vision-language models are trained to describe images with fluent, contextually appropriate sentences. However, these conventional systems are designed for fixed, closed-world scenarios in which all training data are available upfront, and the models are not required to adapt to new visual categories or linguistic concepts over time. In contrast, continual learning is an emerging research paradigm that seeks to endow captioning systems with the ability to learn from a stream of data composed of sequentially arriving tasks, new domains, or novel object categories, all without revisiting previous data [2, 3]. This setting introduces unique challenges associated with lifelong learning, such as the risk of catastrophic forgetting [4], in which previously acquired knowledge is overwritten as the model adapts to novel information. This paradigm can be applied to the case of image captioning, [5, 6]. Successful continual image captioning systems would be capable of incrementally expanding their understanding and generating accurate, coherent captions for both seen and unseen concepts as they are exposed to them. The motivation for continual image captioning extends beyond academic interest. It is highly relevant for real-world applications such as autonomous agents, assistive devices, and interactive robotics, which operate in dynamic environments and must continually assimilate new visual and linguistic patterns [7]. Unlike classification, captioning is a generative task that demands cross-modal reasoning—grounding visual information in language with both fine-grained perceptual detail and pragmatic contextualization. As such, continual image captioning provides a rich testbed for studying not only incremental learning and memory-efficient architectures, but also the complex interplay between vision and language adaptation [8]. It invites exploration of new training strategies, self-supervised learning, and regularization schemes specifically tailored for vision-language models in an open world. Despite growing interest, continual learning methods for computer vision have been largely explored in the context of image classification [3], with comparatively less attention to more complex vision-language generation settings. Prior attempts to mitigate catastrophic forgetting using prompt-based methods or contrastive learning have shown preliminary promise in related tasks [9, 10], but unique challenges arise for generative modeling: the need for preserving object-word alignments, maintaining linguistic diversity, and avoiding loss of semantic granularity across tasks. In this work, we address these challenges by introducing a novel multi-loss training framework for continual image captioning. Building on a pretrained ViT-GPT-2 architecture [11], our approach combines the standard cross-entropy loss with (1) a prompt-based cosine similarity loss ($L_{\text{nouns}}$) to align image embeddings with semantic prompt representations, (2) a CLIP-based cosine similarity loss ($L_{\text{CLIP}}$) to align image embeddings to target caption textual embedding. The framework is initiated with prompt-based, later switching to caption-based alignment training. (3) A language-guided contrastive loss on the class level ($L_{\mathrm{LGCL}}$), similar to [12], which promotes semantic discriminability across tasks through a triplet-based cosine similarity scheme. Our proposed dynamic loss balancing mechanism further prevents any single loss component from dominating training, reducing the risk of overfitting or under-adaptation. All four losses are summed together into a single loss, which is used for backpropagation. Through experiments on a continual split of the MS-COCO dataset [6], we show that our method outperforms baseline and state-of-the-art approaches, in particular exhibiting improved semantic retention. Note that our method incurs no inference-time overhead, making it attractive for practitioners in resource-constrained environments. Our contributions can be summarized as follows.
  • We propose a new prompt-based training approach for continual image captioning that reduces catastrophic forgetting as compared to state-of-the-art-methods, by employing a novel composite objective to align image and text embeddings in a continual learning setting.
  • We release code and standardized dataset splits for the two continual MS-COCO benchmarks used in our experiments, enabling objective and reproducible comparisons.
2 Related Work
Image captioning, the task of generating descriptive natural language sentences for images, has seen remarkable advances through deep learning. Early approaches, such as the Neural Image Caption Generator (NIC) [13], demonstrated promising results and inspired a series of works aimed at improving accuracy and expressiveness [14]. Despite progress, these models are typically trained offline on fixed datasets and crucially suffer from catastrophic forgetting when new tasks are encountered sequentially, a phenomenon well documented in the continual learning literature [4], [3]. To address catastrophic forgetting, continual learning methods have emerged and can be largely classified into the following categories of approaches: regularization-based, replay-based, optimization-based, representation-based, architecture-based [3]. Prompt-based continual learning is a recent approach that belongs to the category of representation-based approaches and introduces parameter-efficient solutions by leveraging pre-trained models and learnable prompts, synthesizing transfer learning ideas from natural language processing (NLP) and extending them to vision-language models [9], [10], [12]. This approach bypasses the need to store or replay task data and can encode task information directly in prompts without requiring explicit task IDs [15]. Recent work integrating language guidance into vision models has demonstrated improved generalization, especially for unseen classes, by aligning visual and semantic spaces [16], [17], [18]. Within image captioning, the challenge of continual learning is rarely studied. Methods such as ContCap [6] employed freezing, pseudo-labeling, and feature distillation. The approaches [5], [19] focussed particularly on recurrent approaches. Yet prompt-based approaches in image captioning remain largely unexplored. Our research advances this frontier by integrating prompt-based continual learning and language guidance, establishing a novel, efficient paradigm for incremental image captioning that addresses catastrophic forgetting without rehearsal or extensive retraining.
3 Method
3.1 Notation
Continual Learning focuses on training a machine learning model on a data stream originating from a sequence of tasks. Let us denote this sequence of tasks as $\mathcal{D}=\{D_{1},D_{2},\ldots,D_{T}\}$, where each task $D_{t}$ consists of tuples $(x_{tk},y_{tk})$. Here, $x_{tk}\in\mathcal{X}$ represents the input image $k$ in task $t$ in the RGB color space, and $y_{tk}\in\mathcal{Y}$ denotes the corresponding label associated with task $t$. For the vocabulary we utilize the pretrained GPT-2 tokenizer [20], with a vocabulary of $50,257$ tokens. This vocabulary, originally trained on a WebText corpus, allows for efficient and consistent tokenization of textual inputs without modifying the underlying language model. By leveraging this fixed vocabulary, our method ensures compatibility with the GPT-2 architecture.
3.2 Proposed Approach
To address the challenge of producing image captions that are both semantically rich and visually discriminative and effective against catastrophic forgetting, we introduce a training framework that integrates multiple supervisory signals. The proposed approach, illustrated schematically in Figure 2,
Figure 2: Overview of the proposed multi-objective training approach for image captioning. The model is optimized with four supervisory signals: Standard Cross-Entropy Loss, Prompt based Cosine Similarity Loss, CLIP based Cosine Similarity Loss and Language-Guided Contrastive Loss.
Standard Cross-Entropy Loss ($L_{\mathrm{CE}}$)
The cross-entropy loss is defined as:
$L_{\mathrm{CE}}(\theta)=-\sum_{t=1}^{T}\sum_{i=1}^{V}y_{t}^{i}\log\hat{y}_{t}^{i},$
where $T$ is the number of words in the ground truth caption, $V$ is the vocabulary size, $y_{t}^{i}$ is the binary indicator (one-hot encoding) of the true target at position $t$, $\hat{y}_{t}^{i}=p_{\theta}(y_{t}=i|y_{1:t-1},I)$ is the model-predicted probability for the $i$-th vocabulary term at position $t$, given the image $I$, the model weights $\theta$ and the previous words ($y_{1:t-1}$). This loss encourages the model to generate linguistically coherent and contextually appropriate captions from image encodings.
Prompt-Based Cosine Similarity Loss ($L_{\mathrm{nouns}}$)
To explicitly reinforce the model’s sensitivity to the key visual elements within a scene, we employ a cosine similarity loss between the image representation and a noun, adjective and action centric prompt embeddings. For a given reference caption $Y$, we extract a subset of tokens corresponding to salient visual entities such as objects (nouns), attributes (adjectives), and actions (verbs). These are arranged into textual prompts, following the GPT-2 format, which is then encoded by a frozen GPT-2 model. Denoting the normalized image and prompt embeddings as $\mathbf{e(\theta)}_{\text{img}}$ and $\mathbf{e}_{\text{prompt}}$ respectively. Based on the cosine similarity:
$\cos(\mathbf{u},\mathbf{v})=\frac{\mathbf{u}^{\top}\mathbf{v}}{\|\mathbf{u}\|\cdot\|\mathbf{v}\|}.$
We define the loss as
$L_{\mathrm{nouns}}(\theta)=1-\cos\left(\mathbf{e(\theta)}_{\text{img}},\mathbf{e}_{\text{prompt}}\right).$
This formulation ensures that the visual encoder produces embeddings that align semantically with linguistically grounded representations of key entities in the scene. It acts as an auxiliary signal in early training (for our experiments ($L_{\mathrm{nouns}}$) was first calculated for first 2 epochs) to improve concept grounding and feature localization.
CLIP-Based Cosine Similarity Loss ($L_{\mathrm{CLIP}}$) In later training epochs, we transition from structured prompts to using the model’s own target decoded captions for alignment through a Cosine Similarity loss between the image embedding and the embedding of the target decoded captions. Denoting the normalized image and target caption embeddings as $\mathbf{e}_{\text{img}}$ and $\mathbf{e}_{\text{caption}}$ respectively:
$L_{\mathrm{CLIP}}(\theta)=1-\cos\left(\mathbf{e(\theta)}_{\text{img}},\mathbf{e}_{\text{caption}}\right).$
This loss encourages alignment between visual and textual modalities in a shared semantic space, reinforcing multimodal consistency as the model becomes more capable of producing fluent and context-aware descriptions.
Language-Guided Contrastive Loss ($L_{\mathrm{LGCL}}$) Following recent advances in language guidance for prompt-based continual learning [12], we further enhance the model’s discriminative capacity by employing a language-guided contrastive objective. Here, we draw inspiration from triplet loss formulations [21], aiming to maximize the alignment between semantically corresponding image-text pairs while repelling mismatching associations.
Let $\mathbf{v(\theta)}_{\text{img}}$ be the embedding of the image, $\mathbf{v}_{\text{text}}^{+}$ the embedding of the positive prompt or caption (derived from the current task/noun prompt), and $\mathbf{v}_{\text{text}}^{-}$ the embedding of a negative prompt (e.g., an unrelated caption from a previous task). The triplet contrastive loss is then expressed as:
$L_{\mathrm{LGCL}}(\theta)=1-\cos(\mathbf{v(\theta)}_{\text{img}},\mathbf{v}_{\text{text}}^{+})+\cos(\mathbf{v(\theta)}_{\text{img}},\mathbf{v}_{\text{text}}^{-}).$
This objective encourages $\mathbf{v(\theta)}_{\text{img}}$ to remain proximate to its correct linguistic counterpart, while maintaining a margin from negatives.
Input: Batch of images, input_ids, labels, prompts, task_num, epoch
Output: Updated model parameters minimizing total loss
for (images, input_ids, labels, prompts) in train_dl do
 $\mathcal{L}_{\text{ce}},\mathbf{e}_{\text{img}},\_\leftarrow$ model(images, input_ids, labels)
 $\mathcal{L}_{\text{lgcl}}\leftarrow 0$, $\mathcal{L}_{\text{nouns}}\leftarrow 0$, $\mathcal{L}_{\text{clip}}\leftarrow 0$
 if use_lgcl then
      prompt_ids $\leftarrow$ tokenizer(prompts)
    $\mathbf{e}_{\text{img}}\leftarrow$ normalize($\mathbf{e}_{\text{img}}$)
    if epoch $\leq 2$ then
       $\mathbf{e}_{\text{nouns}}\leftarrow$ normalize(encode_text(prompt_ids))
       $\mathcal{L}_{\text{nouns}}\leftarrow 1$- cosine_similarity($\mathbf{e}_{\text{img}},\mathbf{e}_{\text{nouns}}$)
    else
         decoded $\leftarrow$ decode_captions(labels)
         caption_ids $\leftarrow$ tokenizer(decoded)
       $\mathbf{e}_{\text{caption}}\leftarrow$ normalize(encode_text(caption_ids))
       $\mathcal{L}_{\text{clip}}\leftarrow 1$- cosine_similarity($\mathbf{e}_{\text{img}},\mathbf{e}_{\text{caption}}$)
      end if
    if task_num == 0 then
       $\mathbf{e}_{\text{pos}}\leftarrow$ normalize(encode_text(prompt_ids))
       foreach $\vec{v}$ in $\mathbf{e}_{\text{pos}}$ do
            append $\vec{v}$ to current_task_pool
         end foreach
      end if
    if task_num > 0 then
       $\mathbf{e}_{\text{pos}}\leftarrow$ normalize(encode_text(prompt_ids))
       foreach $\vec{v}$ in $\mathbf{e}_{\text{pos}}$ do
            append $\vec{v}$ to current_task_pool
         end foreach
       if len(neg_prompt_pool) $\geq B$ then
          $\mathbf{E}_{\text{neg}}\leftarrow$ normalize(stack(subset of neg_prompt_pool))
          $\mathbf{S}\leftarrow\mathbf{e}_{\text{img}}\cdot\mathbf{E}_{\text{neg}}^{\top}$
          $j^{*}\leftarrow\arg\min_{j}\mathbf{S}_{ij}$
          $\mathbf{e}_{\text{neg}}^{(i)}\leftarrow\mathbf{E}_{\text{neg}}[j^{*}]$
         end if
       else
          $\mathbf{e}_{\text{neg}}\leftarrow$ normalize(broadcast(neg_prompt_pool[0]))
         end if
       $s^{+}\leftarrow$ cosine_similarity($\mathbf{e}_{\text{img}},\mathbf{e}_{\text{pos}}$)
       $s^{-}\leftarrow$ cosine_similarity($\mathbf{e}_{\text{img}},\mathbf{e}_{\text{neg}}$)
       $\mathcal{L}_{\text{lgcl}}\leftarrow\text{mean}(\max(0,1-s^{+}+s^{-}))$
      end if
   end if
 // Total loss
 $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{ce}}+\mathcal{L}_{\text{lgcl}}+\mathcal{L}_{\text{nouns}}+\mathcal{L}_{\text{clip}}$
 // Backpropagate and optimize:
    $\mathcal{L}_{\text{total}}.\texttt{backward()}$, optimizer.step(), optimizer.zero_grad()
end for
Algorithm 1 Continual Captioning Training with Language-Guided Losses (LGCL, Nouns, CLIP)
Input: Image path, temperature ($t$), deterministic flag
Output: Generated caption, tokens, logits, loss
// Load and preprocess image
image $\leftarrow$ Image.open(image_path).convert(’RGB’)
image $\leftarrow$ Apply transforms (resize, normalize, to tensor)
image $\leftarrow$ image.unsqueeze(0).to(device)
// Initialize sequence with BOS token
sequence $\leftarrow$ tensor([BOS_TOKEN_ID]).to(device)
// Generate caption autoregressively
tokens, logits, loss $\leftarrow$ model.generate( image, sequence, max_tokens=50, temperature=$t$, deterministic=deterministic )
// Decode tokens to text
caption $\leftarrow$ tokenizer.decode(tokens)
return caption, tokens, logits, loss
Algorithm 2 Inference Algorithm (Caption Generation)
4 Experiments
4.1 Metrics
We evaluated the performance of our models using standard image captioning metrics, including BLEU [22], ROUGE [23], CIDEr [24], and METEOR [25], used in COCO-Caption. Note, METEOR tracks semantic consistency particularly well at the sentence level, because it aligns hypotheses to references using stems, synonyms as well as paraphrase tables and balances precision and recall [26]. We also report CLIPScore [8], which evaluates caption quality by measuring the cosine similarity between the CLIP image and text embeddings and correlates very well with human judgment[8]. In all cases, higher scores indicate better performance. To quantify forgetting, we consider an end-of-task forgetting metric and compute the relative change in metric performance for each task as follows:
[…]
This expresses the proportion of performance lost due to forgetting.
4.2 Experimental Setup
All experiments were performed on a system equipped with an NVIDIA GeForce RTX 4060 Ti GPU. For both ContCap and RATT dataset splits, the setup remained the same. We trained the model for $5$ epochs, with the batch size of 32 and a learning rate (LR) of 1e-5. We used the AdamW optimizer[27], with default momentum parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a base learning rate of $\eta_{0}=\text{LR}/25$, where LR is the peak learning rate.
4.3 Comparison to State-of-the-Art Methods
Table 1, in the same tabular format as in the original paper [6] (Table III), shows that our CLICITA model achieves substantial improvements over the best ContCap baselines ($S_{19}$ and $S_{multiple}$) in nearly all key evaluation metrics. The most notable gains are observed in CIDEr, METEOR, and BLEU-4, indicating that CLICITA significantly improves both the relevance and the fluency of generated captions. While ROUGE-L lags slightly behind the best ContCap variant.
Table 1: Comparison of Best ContCap ($S_{19}$ and $S_{multiple}$) vs. CLICITA
Metric Best in ContCap ($S_{19}$) Best in ContCap ($S_{multiple}$) CLICITA Improvement ($S_{19}$) Improvement ($S_{multiple}$)
BLEU-1 47.1 ($FD$) 53.6 ($FD$) 67.10 +20.00 +13.50
BLEU-4 6.6 ($P$) 10.5 ($P$) 22.00 +15.40 +11.50
ROUGE-L 34.0 ($FD$) 40.0 ($FD$) 36.10 +2.10 -3.90
METEOR 11.2 ($D_{F}$) 14.5 ($E_{F}$) 27.90 +16.70 +13.40
CIDEr 10.0 ($P$) 19.2 ($E_{F}$) 66.20 +56.20 +47.00
Table 2 presents a task-wise comparison on the RATT split, in similar tabular format as in the [5], (Table 2). While CLICITA reports lower BLEU-4 and CIDEr scores across all tasks, indicating a slight drop in n-gram precision and content specificity, it consistently outperforms RATT on METEOR, a key semantic metric. This suggests that CLICITA produces captions with better overall meaning alignment and linguistic fluency, even when exact n-gram matches are fewer.
Table 2: Comparison of RATT vs. Our Method (CLICITA) across tasks. All values reported after training on the last task.
Metric Transport Animals Sports Food Interior
RATT CLICITA $\Delta$ RATT CLICITA $\Delta$ RATT CLICITA $\Delta$ RATT CLICITA $\Delta$ RATT CLICITA $\Delta$
BLEU-4 21.26 17.50 -3.76 24.68 19.70 -4.98 31.61 19.90 -11.71 21.69 18.50 -3.19 27.27 19.80 -7.47
METEOR 21.69 26.50 +4.81 23.49 30.30 +6.81 27.07 28.90 +1.83 21.10 28.40 +7.30 22.57 29.90 +7.33
CIDEr 63.49 58.30 -5.19 72.49 64.30 -8.19 80.85 56.60 -24.25 51.95 50.60 -1.35 65.36 51.50 -13.86
In the following sections we evaluated our method on two different datasets in two distinct settings to investigate the improvement of the designed mechanism over the baseline pretrained model. Note that pre-trained models already serve the continual learning setting due to flat minima more often found in pre-trained models [3]. The models are:
  • CLICITA: Combines all loss functions ($L_{\text{CE}}$, $L_{\text{nouns}}$, $L_{\text{CLIP}}$ and $L_{\text{LGCL}}$).
  • Pre-Trained Basemodel: This is a baseline pre-trained image-captioning model without continual-learning mechanisms.
4.4 Experimental Result - ContCap Dataset Split
Following the training approach outlined in ContCap [6], we incrementally introduced five object classes. These five classes (tasks) were sequentially added one by one into our model: person, sports ball, tv, toilet, and bottle, with the class ’bottle’ being the final task trained for the model. The focus is on knowledge retention. Table 3 shows the average forgetting scores for this dataset.
Table 3: Comparison of average total forgetting across ContCap and RATT Splits. Lower values indicate better knowledge retention. Bold highlights indicate best results per metric.
Split BLEU-1 BLEU-4 ROUGE-L METEOR CIDEr CLIP Average Forgetting
Average Forgetting (%)
CLICITA (ContCap) -0.32 -2.42 -0.96 2.66 -4.42 -0.44 -5.9
Pre-Trained Basemodel (ContCap) -1.18 -3.10 -1.68 3.72 -4.06 -0.36 -6.66
As can be observed from Table 3, our CLICITA method achieves overall better knowledge retention than Pre-Trained Basemodel.
4.5 Experimental Results - RATT dataset split
In RATT [5] split they define five distinct tasks (Transport, Animals, Sports, Food, Interior) based on object categories and processes the dataset to ensure non-overlapping image assignments across tasks. The implementation strictly enforces the paper’s reported image counts per task (14,266/3,431/3,431 for Transport, 9,314/2,273/2,273 for Animals, etc.) by first filtering images containing relevant categories, removing duplicates across tasks, and maintaining only images with more or equal to 5 captions and keeping only the first 5 captions. The validation set is further split 50/50 into validation and test sets. We compare our CLICITA method against Pre-Trained Basemodel, again with a focus is on knowledge retention.
Table 4: Comparison of Average and Total Forgetting Across ContCap and RATT Splits. Lower values indicate better knowledge retention. Bold highlights indicate best results per metric.
Split BLEU-1 BLEU-4 ROUGE-L METEOR CIDEr CLIP Average Forgetting
Average Forgetting (%)
CLICITA (RATT) -0.14 -2.50 -1.36 0.24 -3.62 -0.68 -7.06
Pre-Trained Basemodel (RATT) -0.50 -3.60 -2.00 -0.76 -3.76 -0.60 -11.22
Again, as can be observed in Table 4, our CLICITA method achieves better knowledge retention as compared to Pre-Trained Basemodel. In TAble 5 shows qualitative results of the proposed method after different training stages for different tasks. It can be observed that the captioning is reasonable and semantically similar to the target.
Table 5: Qualitative results on RATT dataset split of our CLICITA model
Task / Target Caption After training each task After training task 5 (Interior)
Task 1: Transport Target: “A passenger bus that is driving down the street” [Uncaptioned image] “public transit bus on a city street.” “public transit bus traveling down a city street.”
Task 2: Animals Target: “A number of zebras standing in the dirt near a wall” [Uncaptioned image] “zebra standing next to a group of zebras.” “zebra standing next to a bunch of other zebras.”
Task 3: Sport Target: “A man is holding a surfboard and staring out into the ocean” [Uncaptioned image] “man carrying a surfboard on the beach.” “man standing in the water holding a surfboard.”
Task 4: Food Target: “A woman sells cupcakes with fancy decorations on them” [Uncaptioned image] “woman is holding a cupcake with a sign on it.” “woman is standing in front of a cupcake display.”
5 Discussion
Overall, across both ContCap and RATT dataset splits and as compared against the original state-of-the-art methods, our CLICITA method achieves mostly better knowledge retention, with lower average forgetting, in particular in metrics addressing semantic consistency. Moreover, CLICITA performs overall favorably compared to the pre-trained basemodel, with respect to average forgetting scores of different metrics in the mentioned datasets. [^1] Note, the scores for each task can be reproduced via the code that will be shared.
6 Conclusion
In this work, we introduced CLICITA, a novel continual image captioning framework that integrates prompt-based semantic guidance with multiple image-text alignment losses. By combining cross-entropy with prompt-based cosine similarity, CLIP-style cosine similarity loss alignment, and language-guided contrastive learning, our method effectively mitigates catastrophic forgetting while maintaining strong caption generation performance. Notably, our approach introduces no inference-time overhead, making it suitable for deployment in resource-constrained environments. Experimental results on continual MS-COCO benchmarks demonstrate that CLICITA significantly outperforms existing ContCap baseline in most of the image captioning metrics, while outperforming RATT baseline in semantic METEOR metric. Future work can explore the method in stronger and more robust models, as well as on a bigger continual learning image captioning dataset.
Acknowledgments
The authors have no competing interests to declare that are relevant to the content of this article. This work was supported by the IU incubator project "Interactive Motion Agent for XR" (IMA-XR), with the project number PR03110311.